OpenStudy (anonymous):

Which fraction is equivalent to 9\frac{3}{5}? A. \frac{48}{5} B. \frac{27}{5} C. \frac{32}{5} D. \frac{27}{45}

OpenStudy (alexandervonhumboldt2):

\[9\frac{ 3 }{ 5}=\frac{ 9*5+3 }{ 5 }\]

OpenStudy (alexandervonhumboldt2):

simplify

OpenStudy (anonymous):

48/5.how would I simplyify t

OpenStudy (alexandervonhumboldt2):

48/5 is the simplified form

OpenStudy (anonymous):

o

OpenStudy (alexandervonhumboldt2):

to find an improper fraction multiply the whole number by denominator and then add numerator as numerator for improper fraction. leave the denominator the same.

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (anonymous):

makes sense

OpenStudy (anonymous):

Thanks

OpenStudy (alexandervonhumboldt2):

np